1/1

Conrad Ferdinand Meyer,

Two sails

tradukita de Leland Bryant Ross

Two sails flashing bright on the royal-blue bay! Two sails fill out tight and sail calmly away!

When one in the breezes moves hither and yon, The other with ease as its shadow responds.

If one wants to hasten, the other goes fast, If one asks for quiet, its companion has rest.

Traduko de la Germana poemo “Zwei Segel” de Conrad Ferdinand Meyer (Konrado Ferdinando Majero, ∗1825-10-11 – †1898-11-28) en la Anglan de Leland Bryant Ross (Ros’ Haruo).

Arg-1088-2195 (2014-09-07 12:48:46)

Tiu ĉi poem-angligo troviĝas en http: // www. thehymnsociety. org/ bbs/ index. cgi? noframes; read= 245 .

http://www.poezio.net/version?poem-id=1088&version-id=,2195